OpenStudy (anonymous):

How many liters of 15.0 molar NaOH stock solution will be needed to make 20.0 liters of a 1.2 molar NaOH solution? Show the work used to solve this problem.

OpenStudy (taramgrant0543664):

So concentration= moles/volume So if I want my concentration to be 1.2M and have a volume of 20, so we need moles (I will represent this with an x) 1.2=x/20 24moles=x So then we know we need 24 moles of NaOH in our solution to get what we want and our initial concentration is 15M so now we need volume (I will represent that with a v) 15=24/v v=24/15 v=1.6L of NaOH is needed